A 5 L flask containing N2 at 1 bar and 298k is connected to a 4 L flask containing N2  at 2 bar and 273k. After the gases are allowed to mix keeping both flasks at their original temp, wod will be the pressure and amount of N2 in the 5 L flask assuming ideal gas behaviour???


(Plzzzz explain me the concept of this question that I am missing)

Grade:12

1 Answers

Aman Bansal
12 years ago

Dear Abhijat,

Diffusion of gases will take place until the pressure of both the flasks is the same,

But the "Total "volume and temperature remains the same during the whole process .

Thus, applying ideal gas equation :

pV=nRT

or p1V1/T1=p2V2/T2

=> 5*1/298  + 4*2/273 = p*9/273+298 /2

Best Of luck

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert